Please help me solve:
What is the slope-intercept form of the line passing through the point (-8,-3) and parallelt to the line 5x+6y=-9

Answer by nerdybill(7384)

What is the slope-intercept form of the line passing through the point (-8,-3) and parallelt to the line 5x+6y=-9
First, determine slope of:
5x+6y=-9
by rearranging into "slope-intercept" form of
y = mx + b
.
5x+6y=-9
6y = -5x-9
y = (-5/6)x - (9/6)
y = (-5/6)x - 3/2
slope is (-5/6)
.
plug slope and point (-8, -3) into
"point-slope" form:
y - y1 = m(x - x1)
y - (-3) = (-5/6)(x - (-8))
y + 3 = (-5/6)(x + 8)
y + 3 = (-5/6)x + (-5/6)8
y + 3 = (-5/6)x - 40/6
y + 3 = (-5/6)x - 20/3
y = (-5/6)x - 20/3 - 3
y = (-5/6)x - 20/3 - 9/3
y = (-5/6)x - 29/3 (this is what they're looking for)